From the point inside the square, the distance in three corners are 4,5, and 6, respectively. Find the length of each side of the square.
Answer by jim_thompson5910(35256)

Drawing


Image generated by GeoGebra (free graphing software).

Let x be the side length of the square ABCD with point P inside the square. Negative lengths do not make sense so we'll make x > 0.

Let's make
PA = 4
PB = 5
PC = 6

Which are the given distances from P to each of the three corners.

-----------------------------------------------------------------------------------------

Focus on triangle APB only for now

Let t = angle ABP which is some angle in degrees.

Note: we don't need to know or solve for t itself, but we will use it indirectly through cos(t).

------------------------------------

Using the law of cosines, we can say

c^2 = a^2+b^2-2*a*b*cos(C)

4^2 = x^2+5^2-2*x*5*cos(angle ABP)

16 = x^2+25-10x*cos(t)

------------------------------------

Isolate the expression cos(t)

16 = x^2+25-10x*cos(t)

16-25 = x^2+25-10x*cos(t)-25

-9 = x^2-10x*cos(t)

x^2-10x*cos(t) = -9

x^2-10x*cos(t)-x^2 = -9-x^2

-10x*cos(t) = -9-x^2

10x*cos(t) = x^2+9

cos(t) = (x^2+9)/(10x) ... call this equation (1)

-----------------------------------------------------------------------------------------

Move onto triangle BPC. Focus only on this triangle.

Since angle ABC is a right angle, this makes

angle ABP+angle PBC = 90

t+angle PBC = 90

angle PBC = 90-t

------------------------------------

By the law of cosines,

c^2 = a^2+b^2-2*a*b*cos(C)

6^2 = x^2+5^2-2*x*5*cos(angle PBC)

6^2 = x^2+5^2-2*x*5*cos(90-t)

36 = x^2+25-10x*cos(90-t)

36 = x^2+25-10x * [ cos(90-t) ]

36 = x^2+25-10x * [ cos(90)*cos(t)+sin(90)*sin(t) ]

36 = x^2+25-10x * [ 0*cos(t)+(1)*sin(t) ]

36 = x^2+25-10x*sin(t)

36-25 = x^2+25-10x*sin(t)-25

11 = x^2-10x*sin(t)

x^2-10x*sin(t) = 11

-10x*sin(t) = 11-x^2

sin(t) = (11-x^2)/(-10x)

sin(t) = (x^2-11)/(10x)

-----------------------------------------------------------------------------------------

Now substitute in the pythagorean trig identity

sin(t) = (x^2-11)/(10x)

sqrt(1-cos^2(t)) = (x^2-11)/(10x)

1-cos^2(t) = [(x^2-11)/(10x)]^2

Then plug in cos(t) = (x^2+9)/(10x) which is equation (1) shown above and get everything to one side.

1-cos^2(t) = [(x^2-11)/(10x)]^2

1-(cos(t))^2 = [(x^2-11)/(10x)]^2

1-((x^2+9)/(10x))^2 = ((x^2-11)/(10x))^2

1-((x^2+9)/(10x))^2 - ((x^2-11)/(10x))^2 = 0

-----------------------------------------------------------------------------------------

Let
g(x) = 1-((x^2+9)/(10x))^2 - ((x^2-11)/(10x))^2

solving g(x) = 0, aka finding the roots of g(x), will yield us the potential solutions for x.

This function is a bit messy fourth degree polynomial so let's turn to a graphing calculator.

I'm using GeoGebra to find the roots, but any graphing calculator or software will do.

I get the following approximate solutions:

{x = -7.069593879182, x = -1.421563358924, x = 1.421563358924, x = 7.069593879182}

which is shown visually here


We can toss out x = -7.069593879182 and x = -1.421563358924 since x > 0.

We can also eliminate x = 1.421563358924 because x must be larger than 6 so P is contained inside the square.

The only possible answer is approximately 7.069593879182

-----------------------------------------------------------------------------------------
-----------------------------------------------------------------------------------------

Answer: The side length of the square is approximately 7.069593879182 units.